If Leah is 6 years older than her sister, Sue, & John is 5 years older than Leah, & the total of their ages is 41. Then how old is Sue?: If Leah is 6 years older than her sister, Sue, & John is 5 years older than Leah, & the total of their ages is 41. Then how old is Sue?
Answer by Paul(988)

Leah = 6+x
Sue = x
John = 5+(6+x)
Total sum = 41

6+x+x+5+6+x=41
3x+17=41
3x=24
x=8

8+6=14
14+5=19

Hence, Sue is 8 years old, Leah is 14 years old and John is 19 years old.
Paul.
Question 24162: If Leah is 6 years older than her sister, Sue, & John is 5 years older than Leah, & the total of their ages is 41. Then how old is Sue?: If Leah is 6 years older than her sister, Sue, & John is 5 years older than Leah, & the total of their ages is 41. Then how old is Sue?
Answer by queenofit(120)

Word problems are tricky. look at them carefully. In your question you want to know sue's age. all the other kids have numbers that are added to them, so it is best to let sue's age be the variable x. so let the following be true, and always write it down, it helps you remember.
x= sue's age
x+6= Leah's age
x+6+5= john's age
do you see how I got this? we do not know sue's age so she is x. Leah is 6 years older than Leah so she is x+6. John is 5 years older than Leah who is 6 years older than sue so John is x+6+5.
Your equation is set up like this
sue's age + leah's age + john's age = 41 understand?
x + x+6 + x+6+5 = 41
now add your common factors
3x+17=41 solve for x
3x=41-17
3x=24
x=8
therefore sue must be 8 years old. to check this you can use sue's age of 8
Sue= 8
Leah= 8+6=14 (Sue's age plus 6 years)
John=14+5=19 ( Leah's age plus 5 years)
Sue+Leah+john=41
8+14+19=41
checking my work proved that this is a correct answer for this problem. Always check your work on this type of question. Remember they may ask you for a different person's age, for example John's age, and then the X answer would not be the correct answer.......
Hope this helps